OpenStudy (buttermequeasy):

Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 5x − 2x = 2x + 2 5x + 2 = 2x − 10 The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describes solutions to equations shown.

jhonyy9 (jhonyy9):

5x-2x-2x=2 x=2 5x-2x= -10-2 3x= -12 x= -12/3 x= -4

jhonyy9 (jhonyy9):

the second frase is right